A 1.5 kilowatt heater is used to heat 2.5kg of water from 25 degree celcius to 100 degree celcius in ti seconds. Find the time taken of the temperature of water to be raised.
1
Expert's answer
2018-01-25T12:58:08-0500
P = Q/t, where P – power of heater, Q – heating energy.
Q = Cm(T2-T1) – where C – water heating capacity (4200 J/kg×K)
t = Q/P = (Cm(T2-T1))/P;
t = (4200×2.5×75)/1500 = 525 (s) – time that heater needs to raise the temperature of 2.5 kg of water from 25 degree celcius to 100 degree celcius.